# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and variable-order nonlocal operators

### v2

@@ 2.1. Properties of function spaces
Assuming the contrary, by the Hahn-Banach theorem (e.g. [51, Theorem 5.19]), there exists a nonzero element $h\in L_{q^{\prime}}(\Omega,\mathbb{F}_{0},\mathbb{P};L_{p^{\prime}}(\mathbb{R}^{d}))$ for which
@@ 2.2. Semigroups and generators of Subordinate Brownian Motions
Characteristic exponent of SBM. It is well known (e.g. [55, Theorem 3.2]) that Bernstein function $\phi$ has a unique representation
@@
We say that $b_{\phi}$ is the drift and the nonnegative measure $m_{\phi}$ satisfying (2.3) is the Lévy measure of $\phi$. According to [54, Theorem 30.1], the SBM $X$ is a $\mathbb{R}^{d}$-valued Lévy process, and its Lévy triplet $(a,D,\nu)$ is as follows:
@@
$b_{\phi}$ and $m_{\phi}$ are the drift and the Lévy measure of $\phi$, respectively. By making use of [54, Theorem 8.1], we get
@@
$\int_{\mathbb{R}^{d}}(\mathrm{e}^{\mathrm{i}\xi\cdot y}-1-\mathrm{i}y\cdot\xi 1_{|y|\leq 1})J(y)\mathrm{d}y$ (2.6)
$=\int_{(0,\infty)}\int_{\mathbb{R}^{d}}(\mathrm{e}^{\mathrm{i}\xi\cdot y}-1-\mathrm{i}y\cdot\xi 1_{|y|\leq 1})p(s,y)\mathrm{d}y\,m_{\phi}(\mathrm{d}s)$
@@
According to [54, Theorem 31.5], if $X$ is a SBM with characteristic exponent $\phi$, then $\mathcal{L}_{X}$ is equal to $\phi(\Delta)$ and has the following integro-differential operator representation:
$\mathcal{L}_{X}f(x)=\phi(\Delta)f(x)=b_{\phi}\Delta f+\int_{\mathbb{R}^{d}}\left(f(x+y)-f(x)-\nabla f(x)\cdot y\mathbf{1}_{|y|\leq 1}\right)J(y)\mathrm{d}y.$
@@ 2.3. Functional calculus
[…]
@@
    the operator $-\phi(\Delta)$ has a bounded $H^{\infty}$-calculus on $L_{p}(\mathbb{R}^{d})$ of angle 00.
@@
$(ii)$ By [14, Theorem 11.9], $-\phi(\Delta)$ has bounded imaginary powers. For $f\in C_{c}^{\infty}([0,T]\times\mathbb{R}^{d})$, Lemma 2.5 provides that $u(t,x):=\int_{0}^{t}T_{t-s}f(s,x)\mathrm{d}s$ fulfills
@@
This aligns with the conditions of [49, Theorem 8.7], implying the boundedness of $\mathcal{T}_{2}$ by applying [49, Theorem 8.7] and a density argument.
@@
We remark that Theorem 2.8-($iii$) does not hold when $q=2$ and $p\in(2,\infty)$. For more detail, see [48, Section 6].
2.4. Solvability of semilinear stochastic integro-diffrential equations
@@
Let $\tau\leq T$ be a bounded stopping time, $\gamma\in\mathbb{R}$, and $(p,q)=(2,2)$ or $(p,q)\in[2,\infty)\times(2,\infty)$.
@@
$\|u\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(\tau)}\leq N\left(\|u_{0}\|_{U_{p,q}^{\phi,\gamma+2}}+\|f(0)\|_{\mathbb{H}_{p,q}^{\phi,\gamma}(\tau)}+\|g(0)\|_{\mathbb{H}_{p,q}^{\phi,\gamma+1}(\tau,\ell_{2})}\right),$
where $N=N(N_{0},N_{1},N_{2})$. Here $N_{0},N_{1},N_{2}$ are constants in Theorem 2.8.
@@
By putting $\tilde{u}:=u1_{\text{$[$\kern-2.66003pt$[$}0,\tau\text{$]$\kern-1.04996pt$]$}}$, it suffices to prove for $\tau\equiv T$. Let $\tilde{u}\in\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)$. By Case 1, there exists a unique solution $v\in\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)$ such that
@@
$\leq\theta^{m}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(t)}^{q}+\sum_{k=1}^{m}\binom{m}{k}\theta^{m-k}N^{k}\int_{0}^{t}\frac{(t-s)^{k-1}}{(k-1)!}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(s)}^{q}\mathrm{d}s.$
By taking sufficiently small $\varepsilon>0$, we have
$\|\mathcal{R}^{m}u-\mathcal{R}^{m}v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}\leq\frac{1}{2}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}.$ (2.16)
@@
This section dedicates our efforts to the proof of Theorem 1.15. Unless explicitly mentioned otherwise, we operate under the assumption that Assumptions 1.11, 1.5, 1.13, and 1.14($\delta_{0}$,$\gamma$) are satisfied throughout. Our strategy for demonstrating Theorem 1.15 involves initially establishing a local (in time) solution. We then extend this to a global (in time) solution by leveraging the established existence, uniqueness, and regularity properties of local solutions to infer analogous properties for global solutions. Here are the results for local solutions and the proofs will be given in Sections 3.1, 3.2, 3.3, and 3.4.
@@
Suppose that $(p,q)=(2,2)$ or $(p,q)\in[2,\infty)\times(2,\infty)$, $\gamma\in(0,1)$, and $u_{0}\in U_{p,q}^{\phi,\gamma}$ is nonnegative.
@@
    on  |($0,\tau\text{$]$\kern-1.49994pt$]$}\times\mathbb{R}^{d}$ with initial data $u_{0}$ has a unique solution $u_{m}\in\mathcal{H}_{p,q}^{\phi,\gamma}(\tau)$, where
@@
    (Positivity of local solutions) The solution $u_{m}$ in $(i)$ is nonnegative.
@@
The proof of Theorem 3.1 will be provided in Sections 3.1, 3.2, 3.3 and 3.4. Now, we prove Theorem 1.15.
@@
Referencing Corollary 3.6, we find that
@@
Step 2. Existence. For the proof of existence, let $u_{m}$ denote a solution to (3.1), as described in Theorem 3.1. Given that $u_{m}\in\mathcal{H}_{p,q}^{\phi,\gamma}(T)$ for any finite $T$, Corollary 3.6 ensures
@@
$\sup_{s\leq t}\sup_{x\in\mathbb{R}^{d}}|u_{R}(\omega,s,x)|=\sup_{s\leq t}\sup_{x\in\mathbb{R}^{d}}|u_{m}(\omega,t,x)|\leq R\quad\forall t\leq\tau_{m}^{R}.$
@@
Consequently, we specify the solution $u(t,x)$ in the interval $[0,\tau_{m})$ as
@@
$(ii)$ Consider $\tau_{m}$ as defined in (3.6) to be a stopping time. For any bounded stopping time $\tau$, we have $u=u_{m}\in\mathcal{H}_{p,q}^{\phi,\gamma}(\tau_{m}\wedge\tau)$. Invoking Theorem 3.1-($iii$), we obtain
@@
Let $\tau\leq T$ be a bounded stopping time and $\gamma\in(0,1)$. Suppose that Assumptions 1.11 1.5 $(\delta_{0},\gamma,r)$ and 1.14($\delta_{0},\gamma$) hold. Suppose also that
@@
where $N=N(K,K^{\prime},N_{0},N_{1},N_{2},\nu_{\delta_{0},\gamma,1})$.
@@
The assertion of Theorem 3.1-(i) is derived from a corollary of Theorem 2.10 and is supported by [17, Theorem 2.4.6]. Our aim is to establish the following estimate:
@@
where $\bm{\pi}=(\pi*e_{1},\pi*e_{2},\ldots)$. Assuming the condition (1.17) holds and for $\gamma\leq 0$, we have
@@
Employing Young’s convolution inequality, (1.4), and given $\gamma\in(0,1)$, we deduce that
@@
$\leq N(K,K^{\prime})\|R_{\delta_{0}(2-\gamma)}\|_{L_{1}(\mathbb{R}^{d})}\|u-v\|_{L_{p}(\mathbb{R}^{d})},$
@@
$\leq N(K,K^{\prime})\|R_{\delta_{0}(2-\gamma)-1}\|_{L_{1}(\mathbb{R}^{d})}\|u-v\|_{L_{p}(\mathbb{R}^{d})}$
$\leq N(K,K^{\prime})\|u-v\|_{L_{p}(\mathbb{R}^{d})}.$
For the inequality
@@
The validity of estimate (3.11) is confirmed by references to [9, Lemma 4.4] and [18, Lemma 3]. By amalgamating (3.8), (3.9), and (3.10), the corollary’s proof is concluded. ∎
@@
3.2. Positivity of local solutions: Proof of Theorem 3.1-(ii)
To establish Theorem 3.1-(ii), which concerns the positivity of local solutions, we employ the maximum principle applicable to semilinear stochastic integro-differential equations represented as
Our approach is guided by the proof of Krylov’s maximum principle as detailed in [40, Theorem 2.5] and [41, Theorem 3.1]. To use the Krylov’s argument, we need the following lemma.
@@
    $\|\phi^{1/2}(\Delta)u\|^{2}_{L_{2}(\mathbb{R}^{d})}=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y<\infty,$ (3.12)
    where $j(x)=j(|x|)$ is defined in (2.4).
@@
    $-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y.$
@@
it suffices only to prove for $u\in\mathcal{S}(\mathbb{R}^{d})$, (3.12) holds. Using Parseval’s identity and definition of $\phi^{1/2}(\Delta)$, we have
@@
$=\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}\left(1-\cos{(\xi\cdot y)}\right)j(|y|)\mathrm{d}y\right)\overline{\mathcal{F}[u](\xi)}\mathcal{F}[u](\xi)\mathrm{d}\xi,$
where the last inequality follows from (2.6). Employing the relation $(1-\cos{z})=|1-\mathrm{e}^{\mathrm{i}z}|^{2}/2$, and applying Plancherel’s theorem again, we achieve
$\int_{\mathbb{R}^{d}}\left|\phi^{1/2}(\Delta)u(x)\right|^{2}\mathrm{d}x$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{1}{2}\left|1-\mathrm{e}^{\mathrm{i}\xi\cdot y}\right|^{2}\overline{\mathcal{F}[u](\xi)}\mathcal{F}[u](\xi)j(|y|)\mathrm{d}\xi\mathrm{d}y$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(1-\mathrm{e}^{-\mathrm{i}\xi\cdot y}\right)\overline{\mathcal{F}[u](\xi)}\left(1-\mathrm{e}^{\mathrm{i}\xi\cdot y}\right)\mathcal{F}[u](\xi)j(|y|)\mathrm{d}\xi\mathrm{d}y$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(x-y)\right|^{2}j(|y|)\mathrm{d}x\mathrm{d}y$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y.$
@@
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|r(u(x))-r(u(y))\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y$ (3.14)
$\leq N\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y=N\|\phi^{1/2}(\Delta)u\|_{L_{2}(\mathbb{R}^{d})}^{2}<\infty.$
@@
$=-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u_{n}(x)-u_{n}(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y.$
@@
$\left|\mathcal{I}(u_{n},v)+\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y\right|$
@@
Due to Lemma 3.3-($iii$), for $u,v\in H_{2}^{\phi,1}(\mathbb{R}^{d})$, the term

### v3

@@ 2.1. Properties of function spaces
Assuming the contrary, by the Hahn-Banach theorem (e.g. [52, Theorem 5.19]), there exists a nonzero element $h\in L_{q^{\prime}}(\Omega,\mathbb{F}_{0},\mathbb{P};L_{p^{\prime}}(\mathbb{R}^{d}))$ for which
@@ 2.2. Semigroups and generators of Subordinate Brownian Motions
Characteristic exponent of SBM. It is well known (e.g. [57, Theorem 3.2]) that Bernstein function $\phi$ has a unique representation
@@
We say that $b_{\phi}$ is the drift and the nonnegative measure $m_{\phi}$ satisfying (2.3) is the Lévy measure of $\phi$. According to [56, Theorem 30.1], the SBM $X$ is a $\mathbb{R}^{d}$-valued Lévy process, and its Lévy triplet $(a,D,\nu)$ is as follows:
@@
$b_{\phi}$ and $m_{\phi}$ are the drift and the Lévy measure of $\phi$, respectively. By making use of [56, Theorem 8.1], we get
@@
$\int_{\mathbb{R}^{d}}(\mathrm{e}^{\mathrm{i}\xi\cdot y}-1-\mathrm{i}y\cdot\xi 1_{|y|\leq 1})j(|y|)\mathrm{d}y$ (2.6)
$=\int_{\mathbb{R}^{d}}(\cos(\xi\cdot y)-1)j(|y|)\mathrm{d}y$
$=\int_{(0,\infty)}\int_{\mathbb{R}^{d}}(\mathrm{e}^{\mathrm{i}\xi\cdot y}-1-\mathrm{i}y\cdot\xi 1_{|y|\leq 1})p(s,|y|)\mathrm{d}y\,m_{\phi}(\mathrm{d}s)$
@@
According to [56, Theorem 31.5], if $X$ is a SBM with characteristic exponent $\phi$, then $\mathcal{L}_{X}$ is equal to $\phi(\Delta)$ and has the following integro-differential operator representation:
$\mathcal{L}_{X}f(x)=\phi(\Delta)f(x)=b_{\phi}\Delta f+\int_{\mathbb{R}^{d}}\left(f(x+y)-f(x)-\nabla f(x)\cdot y1_{|y|\leq 1}\right)J(y)\mathrm{d}y.$
@@ 2.3. Functional calculus
    $\eta_{H^{\infty}}(A):=\inf\{\sigma\in(\eta(A),\pi):A\text{ has a bounded $H^{\infty}(\Sigma_{\sigma})$-calculus}\},$
@@
    the operator $\phi(\Delta)$ has a bounded $H^{\infty}$-calculus on $L_{p}(\mathbb{R}^{d})$ of angle 0.
@@
$(ii)$ By [14, Theorem 11.9], $\phi(\Delta)$ has bounded imaginary powers. For $f\in C_{c}^{\infty}([0,T]\times\mathbb{R}^{d})$, Lemma 2.5 provides that $u(t,x):=\int_{0}^{t}T_{t-s}f(s,x)\mathrm{d}s$ fulfills
@@
This aligns with the conditions of [50, Theorem 8.7], implying the boundedness of $\mathcal{T}_{2}$ by applying [50, Theorem 8.7] and a density argument.
@@
We remark that Theorem 2.8-(iiiiii) does not hold when $q=2$ and $p\in(2,\infty)$. For more detail, see [48, Section 6].
2.4. Solvability of semilinear stochastic integro-differential equations
@@
Let $\tau\leq T$ be a bounded stopping time, $\gamma\in\mathbb{R}$, and $p,q\in(2,\infty)$. Suppose that $f(u)$ and $g(u)$ satisfy the followings:
@@
$\|u\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(\tau)}\leq N\left(\|u_{0}\|_{U_{p,q}^{\phi,\gamma+2}}+\|f(0)\|_{\mathbb{H}_{p,q}^{\phi,\gamma}(\tau)}+\|g(0)\|_{\mathbb{H}_{p,q}^{\phi,\gamma+1}(\tau,\ell_{2})}\right),$ (2.14)
where $N=N(N_{0},N_{1},N_{2},T)$. Here $N_{0},N_{1},N_{2}$ are constants in Theorem 2.8.
@@
By putting
[…]
Let $\tilde{u}\in\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)$. By Case 1, there exists a unique solution $v\in\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)$ such that
@@
$\leq\theta^{m}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(t)}^{q}+\sum_{k=1}^{m}\binom{m}{k}\theta^{m-k}N_{\theta}^{k}\int_{0}^{t}\frac{(t-s)^{k-1}}{(k-1)!}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(s)}^{q}\mathrm{d}s.$
By elementary calculation,
$\sum_{k=1}^{m}\binom{m}{k}\theta^{m-k}N_{\theta}^{k}\int_{0}^{T}\frac{(T-t)^{k-1}}{(k-1)!}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(t)}^{q}\mathrm{d}t$
$\leq\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}\sum_{k=1}^{m}\binom{m}{k}\theta^{m-k}N_{\theta}^{k}\int_{0}^{T}\frac{(T-t)^{k-1}}{(k-1)!}\mathrm{d}t$
$=\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}\sum_{k=1}^{m}\frac{\binom{m}{k}\theta^{m-k}(TN_{\theta})^{k}}{k!}$
$\leq\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}2^{m}\theta^{m}\sum_{k=1}^{\infty}\frac{(TN_{\theta}/\theta)^{k}}{k!}=\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}2^{m}\theta^{m}\mathrm{e}^{TN_{\theta}/\theta}.$
If we take $\theta=1/8$ and $m\in\mathbb{N}$ such that $\mathrm{e}^{TN_{\theta}/\theta}<2^{m}$, then we have
$\|\mathcal{R}^{m}u-\mathcal{R}^{m}v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}\leq\frac{1}{2^{m}}\|u-v\|_{\mathcal{H}_{p,q}^{\phi,\gamma+2}(T)}^{q}.$ (2.17)
@@
This section dedicates our efforts to the proof of Theorem 1.17. Unless explicitly mentioned otherwise, we operate under the assumption that Assumptions 1.5 ($\delta_{0},\gamma,r$), 1.12 ($\delta_{0}$), 1.15, and 1.16($\delta_{0}$$\gamma$) are satisfied throughout. Our strategy for demonstrating Theorem 1.17 involves initially establishing a local (in time) solution. We then extend this to a global (in time) solution by leveraging the established existence, uniqueness, and regularity properties of local solutions to infer analogous properties for global solutions. Here are the results for local solutions and the proofs will be given in Sections 3.1, 3.2, 3.3, and 3.4.
@@
Suppose that $p,q\in(2,\infty)$, $\gamma\in(0,1)$, and $u_{0}\in U_{p,q}^{\phi,\gamma}\cap L_{1}(\Omega\times\mathbb{R}^{d})$ is nonnegative.
@@
[…]
@@
    (Nonnegativity of local solutions) The solution $u_{m}$ in $(i)$ is nonnegative.
@@
The proof of Theorem 3.1 will be provided in Sections 3.1, 3.2, 3.3 and 3.4. Assuming Theorem 3.1 holds, we now proceed to prove Theorem 1.17.
@@
Referencing Theorem 3.1-$(iii)$, we find that
@@
Step 2. Existence. For the proof of existence, let $u_{m}$ denote a solution to (3.1), as described in Theorem 3.1. Given that $u_{m}\in\mathcal{H}_{p,q}^{\phi,\gamma}(T)$ for any finite $T$, Theorem 3.1-$(iii)$ ensures
@@
$\sup_{s\leq t}\sup_{x\in\mathbb{R}^{d}}|u_{R}(\omega,s,x)|=\sup_{s\leq t}\sup_{x\in\mathbb{R}^{d}}|u_{m}(\omega,t,x)|\leq R\quad\forall t\leq\tau_{m}^{R},$
@@
Consequently, we specify the solution $u(t,x)$ in the interval $0,\tau_{m})$ as
@@
$(ii)$ Consider $\tau_{m}$ as defined in (3.6) to be a stopping time. For any bounded stopping time $\tau$, we have $u=u_{m}\in\mathcal{H}_{p,q}^{\phi,\gamma}(\tau_{m}\wedge\tau)$. Invoking Theorem 3.1-(iiiiii), we obtain
@@
Let $\tau\leq T$ be a bounded stopping time, $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$ and $r\in[1,\infty)$. Suppose that Assumptions 1.5 $(\delta_{0},\gamma,r)$, 1.12 ($\delta_{0}$) and 1.16($\delta_{0},\gamma$) hold. Suppose also that
@@
where $N=N(K,K^{\prime},N_{0},N_{1},N_{2},\nu_{\delta_{0},\gamma,1},p,q)$.
@@
To prove the assertion, we apply Theorem 2.10. To this end, it suffices to establish the following estimate, which corresponds to the assumption of Theorem 2.10:
@@
where $\bm{\pi}=(\pi*e_{1},\pi*e_{2},\ldots)$.
Assuming the condition (1.19) holds and for $\gamma\leq 0$, we have
@@
Since $\|R_{\beta}\|_{L_{1}(\mathbb{R}^{d})}=1$ for all $\beta>0$ (see (1.4)), by Young’s convolution inequality and $\gamma\in(0,1)$, we obtain
@@
$\leq N(K,K^{\prime},p)\|R_{\delta_{0}(2-\gamma)}\|_{L_{1}(\mathbb{R}^{d})}^{p}\|u-v\|_{L_{p}(\mathbb{R}^{d})}^{p}$
$=N(K,K^{\prime},p)\|u-v\|_{L_{p}(\mathbb{R}^{d})}^{p},$
@@
$\leq N(K,K^{\prime},p)\|R_{\delta_{0}(2-\gamma)-1}\|_{L_{1}(\mathbb{R}^{d})}\|u-v\|_{L_{p}(\mathbb{R}^{d})}$
$\leq N(K,K^{\prime},p)\|u-v\|_{L_{p}(\mathbb{R}^{d})}.$
Here, the condition $\delta_{0}(2-\gamma)>1$ is ensured by Assumption 1.16($\delta_{0},\gamma$). For the inequality
@@
Here, the constant $\nu_{\delta_{0},\gamma,1}$ is given by
$\nu_{\delta_{0},\gamma,1}:=\int_{\mathbb{R}^{d}}R_{2\delta_{0}(1-\gamma)}(x)\,\pi(\mathrm{d}x),$
which is finite by Proposition 1.6. The validity of estimate (3.11) is confirmed by references to [9, Lemma 4.4] and [18, Lemma 3]. By amalgamating (3.8), (3.9), and (3.10), the corollary’s proof is concluded. ∎
@@
3.2. Nonnegativity of local solutions: Proof of Theorem 3.1-(ii)
To establish Theorem 3.1-(ii), which concerns the nonnegativity of local solutions, we employ the maximum principle applicable to semilinear stochastic integro-differential equations represented as
$\mathrm{d}u=(\phi(\Delta)u+f(u))\mathrm{d}t+\sum_{k=1}^{\infty}g^{k}(u)\mathrm{d}w^{k}_{t}.$
Our approach is guided by the proof of Krylov’s maximum principle as detailed in [40, Theorem 2.5] and [41, Theorem 3.1]. To use Krylov’s argument, we need the following lemma.
@@
    $\|\phi^{1/2}(\Delta)u\|^{2}_{L_{2}(\mathbb{R}^{d})}=b_{\phi}\|\nabla u\|_{L^{2}(\mathbb{R}^{d})}^{2}+\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y<\infty,$ (3.12)
    where $b_{\phi}$ is the drift of $\phi$, and $j$ is defined in (2.4).
@@
    $b_{\phi}\int_{\mathbb{R}^{d}}\overline{\nabla v(x)}\nabla u(x)\mathrm{d}x-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y.$
@@
it suffices only to prove for $u\in\mathcal{S}(\mathbb{R}^{d})$, (3.12) holds. Using Plancherel’s theorem and definition of $\phi^{1/2}(\Delta)$, we have
@@
$=\int_{\mathbb{R}^{d}}\left(b_{\phi}|\xi|^{2}+\int_{\mathbb{R}^{d}}\left(1-\cos{(\xi\cdot y)}\right)j(|y|)\mathrm{d}y\right)\overline{\mathcal{F}[u](\xi)}\mathcal{F}[u](\xi)\mathrm{d}\xi,$
where the last equality follows from (2.5), (2.6) and (2.7). Employing the relation $(1-\cos{z})=|1-\mathrm{e}^{\mathrm{i}z}|^{2}/2$, and applying Plancherel’s theorem again, we achieve
$\int_{\mathbb{R}^{d}}\left|\phi^{1/2}(\Delta)u(x)\right|^{2}\mathrm{d}x$
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{1}{2}\left|1-\mathrm{e}^{\mathrm{i}\xi\cdot y}\right|^{2}\overline{\mathcal{F}[u](\xi)}\mathcal{F}[u](\xi)j(|y|)\mathrm{d}\xi\mathrm{d}y+b_{\phi}\int_{\mathbb{R}^{d}}|\xi|^{2}||\mathcal{F}[u](\xi)|^{2}\mathrm{d}\xi$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(1-\mathrm{e}^{-\mathrm{i}\xi\cdot y}\right)\overline{\mathcal{F}[u](\xi)}\left(1-\mathrm{e}^{\mathrm{i}\xi\cdot y}\right)\mathcal{F}[u](\xi)j(|y|)\mathrm{d}\xi\mathrm{d}y+b_{\phi}\int_{\mathbb{R}^{d}}|\mathrm{i}\xi\mathcal{F}[u](\xi)|^{2}\mathrm{d}\xi$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(x-y)\right|^{2}j(|y|)\mathrm{d}x\mathrm{d}y+b_{\phi}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{2}\mathrm{d}x$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y+b_{\phi}\|\nabla u\|_{L_{2}(\mathbb{R}^{d})}^{2}.$
@@
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|r(u(x))-r(u(y))\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y+b_{\phi}\|\nabla(r(u))\|_{L_{2}(\mathbb{R}^{d})}^{2}$
$\leq N\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|u(x)-u(y)\right|^{2}j(|x-y|)\mathrm{d}x\mathrm{d}y+N\|\nabla u\|_{L_{2}(\mathbb{R}^{d})}^{2}$
$=N\|\phi^{1/2}(\Delta)u\|_{L_{2}(\mathbb{R}^{d})}^{2}<\infty.$
@@
$=$ $-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u_{n}(x)-u_{n}(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y$
$-b_{\phi}\int_{\mathbb{R}^{d}}\overline{\nabla v(x)}\nabla u_{n}(x)\mathrm{d}x.$
@@
$\left|\mathcal{I}(u_{n},v)+\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\overline{v(x)}-\overline{v(y)}\right)\left(u(x)-u(y)\right)j(|x-y|)\mathrm{d}x\mathrm{d}y+b_{\phi}\int_{\mathbb{R}^{d}}\overline{\nabla v(x)}\cdot\nabla u(x)\mathrm{d}x\right|$
@@
Due to Lemma 3.3-(iiiiii), for $u,v\in H_{2}^{\phi,1}(\mathbb{R}^{d})$, the term